Two equally charged particles, held 3.0 x 10-3 m apart, are released from rest. The initial acceleration of the first particle is observed to be 7.9 m/s2 and that of the second to be 9.7 m/s2. If the mass of the first particle is 5.3 x 10-7 kg, what are (a) the mass of the second particle and (b) the magnitude of the charge (in C) of each particle?
Answer #1

Given :-

r = 3 × 10^-3 m

a1 = 7.9 m/s^2

a2 = 9.7 m/s^2

m1 = 5.3 x 10^-7 kg.

m2 = ?

F=ma

Force is equal and opposite

m1a1=m2a2

m2=m1a1/a2

m2 = (5.3 x 10^-7 x 7.9)/9.7

m2 = 4.31 x 10^-7 kg

F = 4.1 x 10^-6 N

F=kq^2/r^2
q = sqrt(Fr^2/k)

q = sqrt(4.1 x 10^-6 x 9 x 10-6)/(9 x 10^9)

q = 6.75 x 10^-11 Coulombs